# Hints-In-Browser: Benchmarking Language Models for Programming Feedback Generation

### v1

@@ Abstract
Generative AI and large language models hold great promise in enhancing programming education by generating individualized feedback and hints for learners. Recent works have primarily focused on improving the quality of generated feedback to achieve human tutors’ quality. While quality is an important performance criterion, it is not the only criterion to optimize for real-world educational deployments. In this paper, we benchmark language models for programming feedback generation across several performance criteria, including quality, cost, time, and data privacy. The key idea is to leverage recent advances in the new paradigm of in-browser inference that allow running these models directly in the browser, thereby providing direct benefits across cost and data privacy. To boost the feedback quality of small models compatible with in-browser inference engines, we develop a fine-tuning pipeline based on GPT-4 generated synthetic data. We showcase the efficacy of fine-tuned Llama3-8B and Phi3-3.8B 4-bit quantized models using WebLLM’s in-browser inference engine on three different Python programming datasets. We will release the full implementation along with a web app and datasets to facilitate further research on in-browser language models.
@@ 1 Introduction
(a) Technique using OpenAI’s GPT-4.
(b) Technique using open-access models.
@@
Figure 1: Different workflows to deploy programming feedback techniques; see Section 1 for details.
@@
Summary of Results. Compared to existing workflows discussed earlier (see Figure 1), hints-in-browser provide clear benefits across two of our new performance criteria: (i) no running costs for an educator or learner; (ii) complete data privacy as the learner’s data does not leave their machine. In fact, after an initial download and caching of the model inside the browser, the learner can continue to benefit from hints-in-browser even without internet connectivity. Given that in-browser inference engines are still in their infancy, they do put stringent requirements on feasible model sizes and consumer hardware. To further enhance and demonstrate the utility of hints-in-browser, we take the following steps:
  • We develop a fine-tuning pipeline using GPT-4 generated synthetic data to boost the feedback quality of small models compatible with in-browser engines. We showcase the efficacy of fine-tuned Llama-3-8B [14, 15] and Phi-3-3.8B [18] 4-bit quantized models using the WebLLM engine [19] on different Python programming datasets and highlight competitive quality surpassing GPT-3.5 [20].
  • We build a web app to demonstrate and reliably measure the performance of hints-in-browser in terms of inference times and functionality across different hardware configurations. We highlight that inference times are competitive with existing feedback generation workflows for compatible machines (e.g., a GPU-equipped consumer laptop, starting at the $1200$ USD price range).
  • We will release the full implementation of our work along with the web app and datasets to facilitate further research and development on hints-in-browser.
@@ 2 Related Work
Small Models and In-Browser Inference. There have been increasing efforts to democratize generative AI technology. On the one hand, this has led to a surge of increasingly powerful small open-access models like Llama-3-8B [14, 15] and Phi-3-3.8B [18], thereby reducing the inference-time compute requirements. On the other hand, this has led to new in-browser inference engines such as WebLLM and ONNX Runtime that enable inference directly in browser [19, 21], without even requiring any installation as needed by end-user technologies like Ollama [22]. We contribute to this ecosystem by benchmarking feedback generation techniques across important real-world performance criteria.
Deployment of Feedback Generation Techniques. There has been an increasing interest in deploying feedback-generation techniques for learners/teachers in real-world settings. Recent research works have demonstrated pilot deployment of programming feedback techniques in classrooms based on OpenAI’s GPT-4 models  [10, 11]. However, given the running costs involved, it is infeasible for educators to provide large-scale free access to such technology. Beyond research, educational websites have also been integrating generative feedback techniques: Khanamigo by Khan Academy [23] and Q-Chat by Quizlet [24] are AI-powered systems based on OpenAI’s GPT-4 models. However, these systems are typically subscription-based to cover the running costs. These recent developments and limitations motivate our work to study and benchmark new deployment workflows.
@@ 3.1 Programming Feedback Scenarios and Technique
Feedback Generation Technique. Figure 2 illustrates the feedback generation technique we consider in this paper adapted from existing techniques [5, 9]. We provide an overview of the key steps and hyperparameters, which inspire the fine-tuning pipeline in Section 4 and affect our results on inference time in Section 5. The starting point is a learner working on a programming task $T$ who has written a buggy program $P_{b}$ and requests a hint. The first step obtains a small set of failing test cases from the test suite $\Omega_{T}$, denoted as $\omega_{T,P_{b}}$. The second step generates a repaired program using the generative model by prompting it with $(T,P_{b},\omega_{T,P_{b}})$. More specifically, we request the model to generate a candidate repaired program, and independently repeat this process $n_{\text{r}}$ times with temperate $t_{\text{r}}$. Then, from this set of $n_{\text{r}}$ programs, we choose the program that passes the test suite $\Omega_{T}$ and is closest to $P_{b}$ in terms of token-edit distance. [^1] Based on Levenshtein edit distance on strings obtained by tokenizing programs using Pygments library [25]. Henceforth, we refer to this repaired program as $P_{r}$; if no repaired program is found, we define $P_{r}$ to be empty. The next step uses all this information and makes a new query to the model asking to generate a concise hint $H$ along with an auxiliary detailed explanation $X$ as part of Chain-of-Thought reasoning [26]. Finally, the hint $H$ is provided to the learner as feedback. [^2] The prompts are adapted from existing literature [9, 11] and will be provided in the supplementary material.
@@ 4 Hints-In-Browser Workflow and Fine-Tuning Pipeline
Hints-In-Browser Workflow. Figure 3 provides an overview of the hints-in-browser deployment workflow and its interaction with a learner requesting hints on a programming task. The starting point is a small language model as a base model in 16-bit format; we will use Llama-3-8B [14, 15] or Phi-3-3.8B [18] throughout this paper. The next stage is to optionally fine-tune the base model to boost the feedback generation quality using the pipeline discussed below. Next, the model is converted to its web version, which also quantizes the model to a 4-bit format suitable for an in-browser inference engine. In this paper, we use the WebLLM engine [19], which makes use of recently added WebGPU capabilities in several modern browsers to enable fast in-browser inference. Next, the web version of the model and auxiliary inference files are hosted in a publicly accessible space such as on HuggingFace [28] or a Github repository [29]. Finally, the hints-in-browser workflow is deployed and accessed by a learner in the form of a web app, which includes a programming environment along with the option to request hints (see Section 6). When the learner accesses the app for the first time, the model is downloaded and cached locally. On subsequent hint requests, our technique runs in the browser and utilizes WebLLM-powered inference to directly provide feedback to the learner. As discussed in Section 1, hints-in-browser provide direct benefits across the performance criteria of cost and data privacy.
@@
Figure 4: Synthetic training data generation process along with shortened version of prompts used for generation. Details are skipped for brevity; we will provide prompts in the supplementary material.
Fine-Tuning Pipeline to Boost Quality. We use the smallest models from Llama-3 and Phi-3 families for hints-in-browser to keep inference times low. To boost the feedback quality of these small models, we develop a fine-tuning pipeline; see Figure 3. The training process uses standard supervised fine-tuning using LoRA [30] and utilizes GPT-4 generated synthetic training data. Next, we discuss the main steps involved in this synthetic training data generation process; see Figure 4(a). The starting point is a set of programming tasks $\{T\}$ that are used for training; this set corresponds to a (sub)set of tasks in the domain that a learner will be working on. The first step is to obtain buggy programs $\{P_{b}\}$ for each task $T$ along with failing test cases $\omega_{T,P_{b}}$ (see Section 3.1 for this notation). These buggy programs can be real-world learner submissions or synthetic programs obtained by prompting GPT-4 – we will consider both settings in the experiments. When using GPT-4 to generate buggy programs, we prompt it for types of real-world mistakes (e.g., syntactical errors, incorrect loop range); see Figure 4(b). The second step is to prompt the GPT-4 model with $(T,P_{b},\omega_{T,P_{b}})$ and ask it to generate a repaired program $P_{r}$, a detailed explanation $X$, and a hint $H$ in a single prompt shown in Figure 4(c); here, we adapt the prompts used in the feedback generation technique. From $(T,P_{b},\omega_{T,P_{b}},P_{r},X,H)$, we assemble four types of training data instances as shown in Figure 4(a). We will provide further implementation details of our synthetic training data generation process in the supplementary material.
@@ 5.1 Programming Domains and Datasets
We consider three domains/datasets representing variety of Python programming tasks; see Figure 5. We provide a high-level overview here and will give the full details in the supplementary material.
@@
BasicAlgo [5, 9]. The first dataset, BasicAlgo, has been introduced in the literature to benchmark generative models and human tutors in terms of feedback generation quality [5, 9]. It comprises $5$ popular Python programming tasks that involve writing basic algorithms, namely GCD, Fibonacci, DivisorsDiv3, Palindrome, and MergeStrs. For each task, the dataset contains $5$ buggy programs based on bugs in real-world submissions on the geeksforgeeks.org platform [31]. These buggy programs represent a variety of non-trivial bugs, including algorithmic mistakes, time/space complexity issues, and Python-specific errors like the mutability of lists. We use these $25$ buggy programs for evaluation. For the fine-tuning pipeline, we begin by generating synthetic buggy programs using GPT-4 in Step-1 of the synthetic training data generation process in Figure 4(a). Then, for each buggy program in the training set, we use GPT-4 to obtain up to $5$ independent tuples $(P_{r},X,H)$; here, we filter invalid tuples, e.g., where $P_{r}$ is not a correct repair. Finally, we assemble $4$ training instances per tuple, which leads to a total of $3304$ instances as reported in Figure 5(a).
@@
KarelAlgo. The third dataset, KarelAlgo, has been designed specifically for this work and involves writing Python programs in a new domain inspired by the Karel visual programming environment [34, 35]. Importantly, these tasks require using domain-specific operations that would have not been seen by the generative models during training. In short, we consider 1-D Karel-worlds which leads to solving linked-list variant of problems, with environment-specific operations (see Figure 5(d)). This dataset comprises $7$ programming tasks that involve writing functions with environment-specific operations, namely Move, CopyEndToAll, SumOfList, NthElementEndOfList, CommonDivisor, Palindrome, and FibonacciSeriesUptoN. For each task, the dataset contains $3$ buggy programs that we designed to capture a mixture of environment-specific bugs and typical bugs in similar problems. We use these $21$ buggy programs for evaluation. For the fine-tuning pipeline, we begin by generating synthetic buggy programs using GPT-4 in Step-1 of the synthetic training data generation process in Figure 4(a) and follow the same process as used for BasicAlgo, which leads to a total of $3676$ instances.
5.2 Models Studied and Quality Assessment
Models Studied. Next, we summarize different models and variants used in the feedback generation technique. Unless specified, we use the same hyperparameters across all models given by the number of repairs $n_{r}=10$, repair temperature $t_{r}=0.7$, and hint temperature of $t_{h}=0.1$. We will provide implementation details about the models and fine-tuning pipeline in the supplementary material.
  • OpenAI’s GPT family. We consider models from OpenAI’s GPT family, including GPT-4 (model=gpt-4-turbo-2024-04-09[1, 20] and GPT-3.5 (model=gpt-3.5-turbo-0125[20]. These models are popularly used in current deployments of programming feedback techniques [10, 11].
@@ 5.2 Models Studied and Quality Assessment
$n_{r}=10$ HGood RPass REdit HGood RPass REdit HGood RPass REdit HGood
@@
Figure 6: Results for feedback quality performance for technique with different models. Averaged results are reported as mean (stderr) and rounded off for brevity. See details in Section 5.3.
@@
$n_{r}=3$ HGood RPass REdit HGood RPass REdit HGood RPass REdit HGood
@@
Figure 7: Results for feedback quality performance with a lower value of $n_{r}$ (total number of program repairs generated by the model as part of the hint generation process). Choosing a higher value can increase hint quality but also increase the inference time. See details in Sections 5.3 and 6.
Quality Assessment. As discussed in Section 3.2, we use quality metrics from literature and report performance for two scenarios of hints generation and program repair. For the program repair quality metrics, we use test suites to automatically compute correctness and edit tokens. For hint quality performance, we follow the procedure used in literature [5, 9] where we ask human evaluators to assess the quality of generated hints w.r.t. the quality attributes. To begin, we did a small-scale investigation to establish the reliability of the rating criteria, where two human evaluators (comprising one author and one external researcher) independently rated $90$ generated feedback instances; we got the Cohen’s kappa reliability value of $0.644$, which indicates substantial agreement between evaluators [36] and match values as observed in literature [9]. Afterward, given the scale of annotations, we did one human annotation per generated feedback for the final evaluation.

### v2

@@ Abstract
Generative AI and large language models hold great promise in enhancing programming education by generating individualized feedback and hints for learners. Recent works have primarily focused on improving the quality of generated feedback to achieve human tutors’ quality. While quality is an important performance criterion, it is not the only criterion to optimize for real-world educational deployments. In this paper, we benchmark language models for programming feedback generation across several performance criteria, including quality, cost, time, and data privacy. The key idea is to leverage recent advances in the new paradigm of in-browser inference that allow running these models directly in the browser, thereby providing direct benefits across cost and data privacy. To boost the feedback quality of small models compatible with in-browser inference engines, we develop a fine-tuning pipeline based on GPT-4 generated synthetic data. We showcase the efficacy of fine-tuned Llama3-8B and Phi3-3.8B 4-bit quantized models using WebLLM’s in-browser inference engine on three different Python programming datasets. We also release the full implementation along with a web app and datasets to facilitate further research on in-browser language models.
@@ 1 Introduction
(a) Workflow using OpenAI’s GPT-4.
(b) Workflow using open-access models.
@@
Figure 1: Deployment workflows for generating programming feedback; see Section 1 for details.
@@
Summary of Results. Compared to existing workflows discussed earlier (see Figure 1), hints-in-browser provides clear benefits across two of our new performance criteria: (i) no running costs for an educator or learner; (ii) complete data privacy as the learner’s data does not leave their machine. In fact, after an initial download and caching of the model inside the browser, the learner can continue to benefit from hints-in-browser even without internet connectivity. Given that in-browser inference engines are still in their infancy, they do put stringent requirements on feasible model sizes and consumer hardware. To further enhance and demonstrate the utility of hints-in-browser, we take the following steps:
  • We develop a fine-tuning pipeline using GPT-4 generated synthetic data to boost the feedback quality of small models compatible with in-browser engines. [^1] Github repository: https://github.com/machine-teaching-group/neurips2024-hints-in-browser-finetuning. We showcase the efficacy of fine-tuned Llama-3-8B [14, 15] and Phi-3-3.8B [18] 4-bit quantized models using the WebLLM engine [19] on different Python programming datasets and highlight competitive quality surpassing GPT-3.5 [20].
  • We build a web app to demonstrate and reliably measure the performance of hints-in-browser in terms of inference times and functionality across different hardware configurations. [^2] Github repository: https://github.com/machine-teaching-group/neurips2024-hints-in-browser-demo. We highlight that inference times are competitive with existing feedback generation workflows for compatible machines (e.g., a GPU-equipped consumer laptop, starting at the $1200$ USD price range).
  • We release the full implementation of our work along with the web app and datasets to facilitate further research and development on hints-in-browser.
@@ 2 Related Work
Small Models and In-Browser Inference. There have been increasing efforts to democratize generative AI technology. On the one hand, this has led to a surge of increasingly powerful small open-access models like Llama-3-8B [14, 15] and Phi-3-3.8B [18], thereby reducing the inference-time compute requirements. On the other hand, this has led to new in-browser inference engines such as WebLLM and ONNX Runtime that enable inference directly in browser [19, 21], without even requiring any installation as needed by end-user technologies like Ollama [22]. Recent works have also evaluated small open-access models for program repair, highlighting their potential in educational settings [23]. We contribute to this ecosystem by benchmarking feedback generation techniques across important real-world performance criteria.
Deployment of Feedback Generation Techniques. There has been an increasing interest in deploying feedback-generation techniques for learners/teachers in real-world settings. Recent research works have demonstrated pilot deployment of programming feedback techniques in classrooms based on OpenAI’s models [10, 11, 24]. However, given the running costs involved, it is infeasible for educators to provide large-scale free access to such technology. Beyond research, educational websites have also been integrating generative feedback techniques: Khanamigo by Khan Academy [25] and Q-Chat by Quizlet [26] are AI-powered systems based on OpenAI’s GPT-4 models. However, these systems are typically subscription-based to cover the running costs. These recent developments and limitations motivate our work to study and benchmark new deployment workflows.
@@ 3.1 Programming Feedback Scenarios and Technique
Feedback Generation Technique. Figure 2 illustrates the feedback generation technique we consider in this paper adapted from existing techniques [5, 9]. We provide an overview of the key steps and hyperparameters, which inspire the fine-tuning pipeline in Section 4 and affect our results on inference time in Section 5. The starting point is a learner working on a programming task $T$ who has written a buggy program $P_{b}$ and requests a hint. The first step obtains a small set of failing test cases from the test suite $\Omega_{T}$, denoted as $\omega_{T,P_{b}}$. The second step generates a repaired program using the generative model by prompting it with $(T,P_{b},\omega_{T,P_{b}})$. More specifically, we request the model to generate a candidate repaired program, and independently repeat this process $n_{r}$ times with temperature $t_{r}$. Then, from this set of $n_{r}$ programs, we choose the program that passes the test suite $\Omega_{T}$ and is closest to $P_{b}$ in terms of token-edit distance. This is based on the Levenshtein edit distance on strings obtained by tokenizing the programs, as done in previous work [9]. Henceforth, we refer to this repaired program as $P_{r}$; if no repaired program is found, we define $P_{r}$ to be empty. The next step uses all this information and makes a new query to the model asking to generate a concise hint $H$ along with an auxiliary detailed explanation $X$ as part of Chain-of-Thought reasoning [27]. Finally, the hint $H$ is provided to the learner as feedback. The prompts used are adapted from existing literature [9, 11] and are provided in our Github repository.
@@ 4 Hints-In-Browser Workflow and Fine-Tuning Pipeline
Hints-In-Browser Workflow. Figure 3 provides an overview of the hints-in-browser deployment workflow and its interaction with a learner requesting hints on a programming task. The starting point is a small language model as a base model in 16-bit format; we will use Llama-3-8B [14, 15] or Phi-3-3.8B [18] throughout this paper. The next stage is to optionally fine-tune the base model to boost the feedback generation quality using the pipeline discussed below. Next, the model is converted to its web version, which also quantizes the model to a 4-bit format suitable for an in-browser inference engine. In this paper, we use the WebLLM engine [19], which makes use of recently added WebGPU capabilities in several modern browsers to enable fast in-browser inference. Next, the web version of the model and auxiliary inference files are hosted in a publicly accessible space such as on HuggingFace [29]. Finally, the hints-in-browser workflow is deployed and accessed by a learner in the form of a web app, which includes a programming environment along with the option to request hints (see Section 6). When the learner accesses the app for the first time, the model is downloaded and cached locally. On subsequent hint requests, our technique runs in the browser and utilizes WebLLM-powered inference to directly provide feedback to the learner. As discussed in Section 1, hints-in-browser provide direct benefits across the performance criteria of cost and data privacy.
@@
Figure 4: Synthetic training data generation process along with shortened version of prompts used for generation. Details are skipped for brevity and full prompts are provided in our Github repository.
Fine-Tuning Pipeline to Boost Quality. We use the smallest models from Llama-3 and Phi-3 families for hints-in-browser to keep inference times low. To boost the feedback quality of these small models, we develop a fine-tuning pipeline; see Figure 3. The training process involves standard supervised fine-tuning with LoRA [30] and uses GPT-4 generated synthetic training data. Next, we discuss the main steps involved in this synthetic training data generation process; see Figure 4(a). The starting point is a set of programming tasks $\{T\}$ that are used for training; this set corresponds to a (sub)set of tasks in the domain that a learner will be working on. The first step is to obtain buggy programs $\{P_{b}\}$ for each task $T$ along with failing test cases $\omega_{T,P_{b}}$ (see Section 3.1 for this notation). These buggy programs can be real-world learner submissions or synthetic programs obtained by prompting GPT-4 – we will consider both settings in the experiments. When using GPT-4 to generate buggy programs, we prompt it for types of real-world mistakes (e.g., syntactical errors, incorrect loop range); see Figure 4(b). The second step is to prompt the GPT-4 model with $(T,P_{b},\omega_{T,P_{b}})$ and ask it to generate a repaired program $P_{r}$, a detailed explanation $X$, and a hint $H$ in a single prompt shown in Figure 4(c); here, we adapt the prompts used in the feedback generation technique. From $(T,P_{b},\omega_{T,P_{b}},P_{r},X,H)$, we assemble four types of training data instances as shown in Figure 4(a). We provide further implementation details of our synthetic training data generation process in our Github repository.
@@ 5.1 Programming Domains and Datasets
We consider three domains/datasets representing a variety of Python programming tasks; see Figure 5. We provide a high-level overview here and defer the full details to the our Github repository.
@@
BasicAlgo [5, 9]. The first dataset, BasicAlgo, has been introduced in the literature to benchmark generative models and human tutors in terms of feedback generation quality [5, 9]. It comprises $5$ popular Python programming tasks that involve writing basic algorithms, namely GCD, Fibonacci, DivisorsDiv3, Palindrome, and MergeStrs. For each task, the dataset contains $5$ buggy programs based on bugs in real-world submissions on the geeksforgeeks.org platform [31]. These buggy programs represent a variety of non-trivial bugs, including algorithmic mistakes, time/space complexity issues, and Python-specific errors like incorrect string indexing. We use these $25$ buggy programs for evaluation. For the fine-tuning pipeline, we begin by generating synthetic buggy programs using GPT-4 in Step-1 of the synthetic training data generation process in Figure 4(a). Then, for each buggy program in the training set, we use GPT-4 to obtain up to $5$ independent tuples $(P_{r},X,H)$; here, we filter invalid tuples, e.g., where $P_{r}$ is not a correct repair. Finally, we assemble $4$ training instances per tuple, which leads to a total of $3304$ instances as reported in Figure 5(a).
@@
KarelAlgo. The third dataset, KarelAlgo, has been designed by us and involves writing Python programs in a new domain inspired by the Karel visual programming environment [34, 35]. Importantly, these tasks require using domain-specific operations that would have not been seen by the generative models during training. In short, we consider 1-D Karel-worlds which leads to solving linked-list variant of problems, with environment-specific operations (see Figure 5(d)). This dataset comprises $7$ programming tasks that involve writing functions with environment-specific operations, namely Move, CopyEndToAll, SumOfList, NthElementEndOfList, CommonDivisor, Palindrome, and FibonacciSeriesUptoN. For each task, the dataset contains $3$ buggy programs that we designed to capture a mixture of environment-specific bugs and typical bugs in similar problems. We use these $21$ buggy programs for evaluation. For the fine-tuning pipeline, we begin by generating synthetic buggy programs using GPT-4 in Step-1 of the synthetic training data generation process in Figure 4(a) and follow the same process as used for BasicAlgo, which leads to a total of $3676$ instances.
5.2 Models Studied and Experimental Details
Models Studied. Next, we summarize different models and variants used in the feedback generation technique. Unless specified, we use the same hyperparameters across all models given by the number of repairs $n_{r}=10$, repair temperature $t_{r}=0.7$, and hint temperature of $t_{h}=0.1$.
  • OpenAI’s GPT family. We consider models from OpenAI’s GPT family, including GPT-4 (model=gpt-4-turbo-2024-04-09[1, 20], GPT-4o (model=gpt-4o-2024-08-06 ), GPT-4o mini (model=gpt-4o-mini-2024-07-18), and GPT-3.5 (model=gpt-3.5-turbo-0125[20]. These models are popularly used in current deployments of programming feedback techniques [10, 11].
@@ 5.2 Models Studied and Experimental Details
Table 1: Results for feedback quality performance for technique with different models. Averaged results are reported as mean (stderr) and rounded off for brevity. See details in Section 5.3.
@@
$n_{r}=10$ HGoodRPassREdit HGoodRPassREdit HGoodRPassREdit HGood%
@@
Table 2: Results for feedback quality performance with a lower value of $n_{r}$ (total number of program repairs generated by the model as part of the hint generation process). Choosing a higher value can increase hint quality but also increase the inference time. See details in Sections 5.3 and 6.
@@
$n_{r}=3$ HGoodRPassREdit HGoodRPassREdit HGoodRPassREdit HGood%
@@
Quality Assessment. As discussed in Section 3.2, we use quality metrics from literature and report performance for the two scenarios of hint generation and program repair. For the program repair quality metrics, we use test suites to automatically compute correctness and edit tokens. For hint quality performance, we follow the procedure used in literature [5, 9] where we ask human evaluators to assess the quality of generated hints w.r.t. the quality attributes. To begin, we did a small-scale investigation to establish the reliability of the rating criteria, where two human evaluators (comprising one author and one external researcher) independently rated $90$ generated feedback instances; we obtained a Cohen’s kappa reliability value of $0.644$, which indicates substantial agreement between evaluators [36] and matches values observed in the literature [9]. Afterward, given the scale of annotations, we did one human annotation per generated feedback for the final evaluation. In our results, we use HGoodand RPassto denote the percentage of evaluation examples that have HGood equal to 1 and RPass equal to $1$, respectively.
Details About Fine-Tuning Experiments. We use the same hyperparameters across all models and domains. For training, we use standard supervised fine-tuning with LoRA [30]. Full implementation details are available in our Github repository. We conducted both the training and inference of the open-access models on a cluster of machines with Nvidia A-100 GPUs.